The acceleration of a particle in one dimensional motion at time t is given by at t () 6 2 = − . Given that the initial velocity of the particle is 0, find the average velocity of this particle over the first 4 seconds.

Please solve and show work. Thank you so much!
Answer #1

alt 6t - Salt) dt 2 = S6-2) dt 6オー2t +C 2. at o, v-o Co O-OC 342-2t +C v-3f-2+ ナこ (342-2dt Vag= L(s#*- 249. Average velocity